You have $14000 to invest and would like to create a portfolio with an expected return of 9.85 percent. You can invest in Stock K with an expected return of 8.3 percent and stock L with an expected return of 11.9 percent. How much will you invest in stock K?
Answer #1

let weight of stock K = X, stock L= (1-x) 0.083 X +0.119(1-x) = 0.0985 0.083 X - 0.119 X = -0.0205 0.5694 weight of: Stock K